OpenStudy (anonymous):

Chelsea invested $15,000 at 4% interest compounded annually. How much interest will she earn in 4 years?

OpenStudy (anonymous):

Pe^(rt) P=principle r= rate t=# years

OpenStudy (anonymous):

She'll get 17602.66 in the end, so interest wise, it;s 2602.66

OpenStudy (anonymous):

A = P(1+0.04)^4 = 15000(1.04)^4 = 15000(1.17) = 17548 Interest = Amount-Principal

OpenStudy (anonymous):

Pe^rt is continuous compounding, you'll want to use douglas's formula